OpenStudy (anonymous):

lim (x,y)---> (0,0) along the x-axis (x+y)^2/(3x^2+4y^2)

OpenStudy (anonymous):

So, you set y=0 \[\lim_{x \rightarrow 0} \frac{x^2}{3x^2} = \frac{1}{3}\]